To transform + = — into a linear differential equation with constant coefficients, di2 dx x the required substitution is (a) x = sin t (b) x=t4+1 (c) x = log t (d) x = el 



1
Expert's answer
2021-02-24T07:38:27-0500

"\displaystyle
(a)\\ x = \sin{t} \\

\frac{\mathrm{d}x}{\mathrm{d}t} = \cos{t} \\

\frac{\mathrm{d}^2x}{\mathrm{d}t^2} = -\sin{t} \\

\frac{\mathrm{d}^2x}{\mathrm{d}t^2} = -x,\,\, \frac{\mathrm{d}^2x}{\mathrm{d}t^2} + x = 0 \\


(b) \\ x = t^4 + 1 \\

\frac{\mathrm{d}x}{\mathrm{d}t} = 4t^3\\

\frac{\mathrm{d}^2x}{\mathrm{d}t^2} = 12t^2\\

\left(\frac{\mathrm{d}^2x}{\mathrm{d}t^2}\right)^2 = 144t^4\\

\left(\frac{\mathrm{d}^2x}{\mathrm{d}t^2}\right)^2 = 144(x - 1)\\

\left(\frac{\mathrm{d}^2x}{\mathrm{d}t^2}\right)^2 - 144(x - 1) = 0\\


(c) \\ x = \log{t} \\

\frac{\mathrm{d}x}{\mathrm{d}t} = \frac{1}{t} \\

\frac{\mathrm{d}x}{\mathrm{d}t} = e^{-x} \\

\frac{\mathrm{d}x}{\mathrm{d}t} - e^{-x} = 0\\



(d) \\ x = e^t \\

\frac{\mathrm{d}x}{\mathrm{d}t} = e^{t} = x \\

\frac{\mathrm{d}x}{\mathrm{d}t} - x = 0"